Solve the following: a. 2015! b. P[6,2] c. C[9,8]
Answer #1

a) 20! / 15!

20! / 15! = 20*19*18*17*16*15*14*13*12*11*10*9*8*7*6*5*4*3*2*1 / 15*14*13*12*11*10*9*8*7*6*5*4*3*2*1

20! / 15! = 20*19*18*17*16

20! / 15! = 1860480

b) P[ 6, 2 ]

P[ n, r ] = n! / r!

P[ 6, 2 ] = 6! / 2!

P[ 6, 2 ] = 6*5*4*3*2*1 / 2*1

P[ 6, 2 ] = 6*5*4*3

P[ 6, 2 ] = 360

c) C[ 9, 8 ]

C[ n, r ] = n! / ( r! * (n-r)! )

C[ 9, 8 ] = 9! / ( 8! * (9-8)! )

C[ 9, 8 ] = 9! / ( 8! * 1! )

C[ 9, 8 ] = 9*8*7*6*5*4*3*2*1 / ( 8*7*6*5*4*3*2*1*1 )

C[ 9, 8 ] = 9